The magnitude of the force (pi+20j) is 29.Find the:
(a) the possible values of p
(b) the direction of the force correct to the nearest degree.

Answer by Alan3354(69443)

The magnitude of the force (pi+20j) is 29.Find the:
(a) the possible values of p
-----------------
p^2 + 20^2 = 29^2
p^2 = 441
p = -21, 21
=========================
(b) the direction of the force correct to the nearest degree.
Angle with the i axis = atan(20/p)